In Joly’s differential steam calorimeter, 3 g of an ideal gas is contained in a rigid closed sphere at 20°C. The sphere is heated by steam at 100°C and it is found that an extra 0.095 g of steam has condensed into water as the temperature of the gas becomes constant. Calculate the specific heat capacity of the gas in J/g-K. The latent heat of vaporization of water = 540 cal/g.

Grade:11

1 Answers

Deepak Patra
9 years ago
Sol. In Joly’s differential steam calorimeter C base v = m base 2L/m base 1(θ base 2 – θ base 1) M base 2 = Mass of steam condensed = 0.095 g, L = 540 Cal/g = 540 × 4.2 J/g M base 1 = Mass of gas present = 3 g, θ base 1 = 20°C, θ base 2 = 100°C ⇒ Cv = 0.095 * 540 * 4.2/3(100 - 20) = 0.89 = 0.9 J/g – K